# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

### v3

@@ Abstract
We present CD-NGP, a fast and scalable representation for 3D reconstruction and novel view synthesis in dynamic scenes. Inspired by continual learning, our method first segments input videos into multiple chunks, followed by training the model chunk by chunk, and finally, fuses features of the first branch and subsequent branches. Experiments on the prevailing DyNeRF dataset demonstrate that our proposed novel representation reaches a great balance between memory consumption, model size, training speed, and rendering quality. Specifically, our method consumes $85\%$ less training memory ($<14$GB) than offline methods and requires significantly lower streaming bandwidth ($<0.4$MB/frame) than other online alternatives. Furthermore, we conduct experiments on a self-collected extra-long video sequences dataset, which demonstrates the superior scalability of our method over previous state-of-the-art approaches. A free copy of this paper and all supplemental materials are available at https://arxiv.org/abs/2409.05166.
@@
The demonstration of our continual learning pipeline for dynamic scenes. Please note the dimensions of representations for different branches remain the same while the complexities of hash tables are flexible and could be configured to ensure scalability. The spatial hash features could be encoded along voxels [31], planes [17] or both [38].
@@ Introduction
In this paper, we propose continual dynamic neural graphics primitives (CD-NGP), a fast and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos. As illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes, we evenly split the dynamic videos into multiple chunks based on the timestamp and conduct reconstruction with different model branches, where each branch includes MLP regressors, spatial encoding for the spatial coordinates $(x,y,z)$, and temporal encoding for the time stamp $t$. We encode the different chunks by hash tables with different complexities to ensure the model scalability and then continually train the model chunk by chunk to alleviate memory occupation. For each individual chunk, we locate the corresponding model branch, sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids [31], and represent spatial features by spatial hash grids. The spatial features of the first branch and the current branch are fused by element-wise sum as default or concatenation if applicable. The temporal features are encoded by the hash table for time stamp $t$ in the current branch only. Spatial and temporal features are concatenated along with viewing direction $d$ and input to MLP regressors for rendering and optimization.
@@ 2.2 Representation of short dynamic intervals
where the $P_{m}(x,y,z)\in\mathbf{R}^{F}$ denotes the tri-linearly interpolated value from the nearby 8 grid vertices w.r.t. 3D coordinates $(x,y,z)$ at the $l$-th resolution, $L$ is the number of the resolutions, and $F$ is the dimensions of features $P_{m}(x,y,z)$, which are similar to those in Instant-NGP [31].
@@ 2.3 Continual dynamic neural graphics primitives (CD-NGP)
where $\Psi_{0}(\mathbf{x})$ denotes the spatial encoder of the base branch, $\Psi_{k}(\mathbf{x})$ denotes the spatial encoder of the current branch. The view-dependent colors are regressed as Eq. 4 with the color MLP $f_{\theta_{2k}}$.
@@ 3.1 Datasets
DyNeRF dataset [28] is the most prevailing dataset for dynamic view synthesis that provides high-quality multi-camera videos, on which most of the current state-of-the-art methods are evaluated. DyNeRF dataset contains 6 individual cooking scenes with varied illumination, food topological changes, and transient effects like flames. It contains five 300-frame videos and one 1200-frame video, with up to 21 synchronized camera views at the resolution of $2704\times 2028$ at 30 frames per second (FPS).
@@
To further prove the viability of our method on long video sequences, we provide a new dataset containing 5 dynamic scenes at $3840\times 2160$ resolution, the lengths of which are at least 200 seconds. Similar to the DyNeRF dataset [28], scenes in our dataset are recorded by 20 GoPro cameras at 30 FPS. We synchronize the videos by environmental sound, and we put chessboard images in the scene to help calibrate the cameras and estimate the camera poses. Compared with those in the DyNeRF dataset, the dynamics are much harder, including large motions of human bodies and transient effects in the background. We collect the dataset for evaluating the scalability of our method and further research. We have checked the videos to make sure they do not contain controversial or offensive content before using them for research purposes. The faces in the videos may contain identifiable information, but the information will neither be used for any other purposes except for research nor be disclosed to any third party. All videos are collected with the explicit consent of all the participants. The view in the upper center is selected for validation.
@@ 3.2 Implementation details
Our implementation is based on HexPlane [7], Instant-NGP [31], and NGP-pl [25]. All the metrics are measured on a single RTX 3090 GPU if not otherwise specified. Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets. Following [7], we exclude the unsynchronized coffee martini scene and conduct experiments on the others. Following most of the state-of-the-art methods [28, 27, 17, 45, 43] we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset. We use $1280\times 720$ resolution for our dataset to fit into memory. We use the importance sampling scheme in [28] at a batch size of 1024 rays. The parameters are optimized by Adam [23], with $\epsilon=10^{-15},\beta=(0.9,0.96)$. The learning rate is set as 0.001 and scheduled by cosine function following [25]. We set the size of the hash table in the first branch as $2^{19}$ and the size of the hash table in the subsequent branches as $2^{14}$, and we set the structural parameters $(L,F)$ in the spatial hash encoders as $(12,2)$ if not otherwise specified. At the base branch, we initialize the models and train for 18k iterations, and for each of the subsequent branches we train for 3k iterations using the previous branch for initialization. We use LeakyReLU activation in the MLP regressors. For long video sequences, we assign 30 chunks as an episiode. The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes. Strategies for updating the occupancy grids are identical with [31]. To save storage, we use only one shared occupancy grid for all the branches. We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision.
@@
Following MERF [38] for static scenes, we use a smaller hash table for the MERF and the MERF-C than pure voxel hash grid [31] representation in Table 5 to ensure the compactness and scalability of our method. Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the MERF and the MERF-C are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively. The number of the parameters of the MERF representation is thus reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids. Since the hash tables take up the majority of the parameters, the model size is reduced from $21M$ to $7.5M$ as in Table 5. We use $2^{P-2}$-length hash tables for the Plane representation in Table 5 to balance the quality and the model size. The structural configurations $(L,F)$ of the plane representation in Table 5 are set as $(6,4)$ for better results.
@@ 3.3 Results and comparisons
Table 1: Comparisons between our CD-NGP and SOTA methods on the DyNeRF dataset. Memory denotes the maximum host memory required for training. - denotes the metrics unreported in the original publications. denotes the methods measured only on the most challenging flame salmon scene. Best, Second, and Third metrics are in highlight.
@@
Table 1 shows the quantitative comparisons between CD-NGP and other state-of-the-art methods on DyNeRF dataset. We use the officially released metrics. The methods in [28, 34] are trained on 8 V100 GPUs or equally 4 A100 GPUs. The methods in [28, 34, 43] have no official implementations available. We are unable to obtain the missing quality metrics in [17] because of their huge memory occupation in the default resolution. HexPlane [7] uses resolution $(1024,768)$ in their implementation, so the memory occupation is much smaller than other offline methods. We use the released output of [27] for DSSIM and LPIPS metrics. StreamRF [27] and INV [48] are trained per-frame. NeRFplayer [43] is trained offline but streamable upon inference. 4DGS [55] caches the frames onto disk for training to alleviate memory occupation with a trade-off on training speed. The reconstruction quality is measured by PSNR, DSSIM, and LPIPS [57] (by AlexNet [24]) following common practices [28, 7, 45]. To prove the viability of our parameter-isolation-based method, we apply the previous state-of-the-art method HexPlane [7] to continual learning setting with parameter-regularization (HexPlane-PR) and generative replay (HexPlane-GR) as additional baselines.
@@
Table 2 and Fig. 5 shows the quantitative and qualitative comparisons on the meetroom dataset. Fig. 6 is the PSNR curve of the compared methods on the trimming scene. As shown in Table 2, HexPlane-GR and HexPlane-PR fail and produce unsatisfactory results on the meetroom dataset, we thus exclude them from Fig. 6. Our method outperforms the implicit MLP-based online method [48] in terms of quality, speed, and model size significantly. Moreover, our method consumes only a fraction of storage than the explicit voxel-based online baseline [27] while maintaining comparable reconstruction quality as compared in Fig. 5. When compared with offline baselines, our method produces comparable quality to MixVoxels [45] and outperforms HexPlane [7] in terms of quality and speed, and consumes much less memory as in Fig. 5. We use the released configuration of [27] for the meetroom dataset. The full model of [27] results in much worse metrics than those reported in their paper, so we only report the metrics of their pilot model.
@@
Ours $\checkmark$ 26.01 0.039 0.1906 77min 113MB
@@
We also provide quantitative comparisons on our proposed long multi-view video dataset in Table 3 and Fig. 7 to further show the advantage in scalability of our method. We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset. We set the hash table size as $2^{12}$ in the auxiliary branches and use $T_{chunk}=5$ in our model. As shown in Table 3, our method reduces the memory and storage consumption by at least 4×4$\times$4 × with comparable quality to HexPlane [7]. Compared with the results in Table 1, the performance gap in storage consumption is even larger in long video sequences. This is because the whole model shares one base branch and the following auxiliary branches occupy the majority of the parameters. We thus set the hash table size in the auxiliary branches to $2^{12}LF$ and leverage the re-use of the features from the base branch to reduce the model size to $\mathcal{O}(N_{chunk}\cdot 2^{12}LF)$. .
The PSNR curves in Fig. 7 further demonstrate the stability of our method. The compared online methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than that in the end. The problem is properly eliminated by our method through parameter isolation, and our method continuously produce high-quality results for the whole sequence. We also provide comparisons between the rendered results in Fig. 8. The regularization method in Fig. 8(a) reconstructs the TV screen in the background erroneously. The replay method in Fig. 8(b) fails to recover the details. Ours in Fig. 8(f) produces comparable quality to the offline baseline in Fig. 8(c).
Figure 7: PSNR curves of the seminar scene in our long video dataset. The solid lines and dotted lines denote offline and online methods respectively.

### v4

@@ Abstract
Current methodologies for novel view synthesis (NVS) in dynamic scenes encounter significant challenges in harmonizing memory consumption, model complexity, training efficiency, and rendering fidelity. Existing offline techniques, while delivering high-quality results, are often characterized by substantial memory demands and limited scalability. In contrast, online methods grapple with the challenge of balancing rapid convergence with model compactness. To address these issues, we propose continual dynamic neural graphics primitives (CD-NGP). Our approach synergizes features from both temporal and spatial hash encodings to achieve high rendering quality, employs parameter reuse to enhance scalability, and leverages a continual learning framework to mitigate memory overhead. Furthermore, we introduce a novel dataset comprising multi-view, exceptionally long video sequences with substantial rigid and non-rigid motion, thereby substantiating the scalability of our method.
We evaluate the quality, speed and scalability our method on both established public datasets and our exceptionally long video dataset. Notably, our method achieves an $85\%$ reduction in training memory consumption (less than 14GB) compared to offline techniques and significantly lowers streaming bandwidth requirements (less than 0.4MB/frame) relative to other online alternatives. The experimental results on our long video sequences dataset show the superior scalability and quality compared to existing state-of-the-art approaches.
@@
The demonstration of our continual learning pipeline for dynamic scenes. Our method segments input videos into multiple chunks, train the model chunk by chunk and leverage the re-use of parameters from the first branch to ensure high scalability. The spatial hash features could be encoded along voxels, orthogonally projected planes or both.
@@ Introduction
In this paper, we propose continual dynamic neural graphics primitives (CD-NGP), a fast and scalable representation for reconstructing dynamic scenes, which enables dynamic reconstruction in long videos.
The proposed methodology is illustrated in CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes: we partition dynamic videos into multiple segments based on timestamps and perform reconstruction using distinct model branches. Each branch includes MLP regressors, spatial hash encoding for spatial coordinates $(x,y,z)$, and temporal hash encoding for the timestamp $t$. For the first segment, we encode spatial features using a large hash table, and for subsequent segments, we use a smaller hash table. By reusing parameters from the initial branch, we enhance scalability and accelerate convergence. For each segment, we identify the corresponding model branch, sample 3D points $(x,y,z)$ according to frame pixels and cached occupancy grids [31], and represent spatial features using spatial hash grids. The spatial features from the initial branch and the current branch are fused through element-wise summation by default or concatenation if applicable. Temporal features are encoded by the hash table for the timestamp $t$ within the current branch only. These spatial and temporal features, along with the viewing direction $d$, are concatenated and input into MLP regressors for rendering and optimization.
@@ 2.2 Representation of short dynamic intervals
where the $P_{m}(x,y,z)\in\mathbf{R}^{F}$ denotes the tri-linearly interpolated value from the nearby 8 grid vertices w.r.t. 3D coordinates $(x,y,z)$ at the $l$-th resolution, $L$ is the number of the resolutions, and $F$ is the dimensions of features $P_{m}(x,y,z)$, which are similar to those in Instant-NGP [31]. If necessary, Eq. 5 could be adopted to the hybrid (MERF) representation [38], defined as:
@@ 2.3 Continual dynamic neural graphics primitives (CD-NGP)
where $\Psi_{0}(\mathbf{x})$ denotes the spatial encoder of the base branch, $\Psi_{k}(\mathbf{x})$ denotes the spatial encoder of the current branch. The view-dependent colors are regressed as Eq. 4 with the color MLP $f_{\theta_{2k}}$. For simplicity, we denote the models adopting pure voxel encoders in Eq. 5 as CD-NGP, the models adopting the hybrid encoders in Eq. 6 as CD-MERF.
@@ 3.1 Datasets
The DyNeRF dataset [28] is the most prevalent benchmark for dynamic view synthesis, providing high-quality multi-camera videos on which current state-of-the-art methods are primarily evaluated. This dataset includes six unique cooking scenes, each featuring varying illumination, food topological changes, and transient effects such as flames. It consists of five videos, each containing 300 frames, and one extended video with 1200 frames, captured at a resolution of $2704\times 2028$ with up to 21 camera views at 30 frames per second (FPS). Following the approach of HexPlane [7], we excluded the unsynchronized scene coffee martini and conducted experiments on the remaining scenes, using the first 300 frames of the 1200-frame video flame salmon for evaluation.
@@
To further validate the effectiveness of our method on long video sequences, we introduce a new dataset comprising 5 dynamic scenes at a resolution of $3840\times 2160$. Similar to the DyNeRF dataset [28], these scenes are captured by 20 GoPro cameras operating at 30 FPS. We synchronize the multi-view videos using environmental sound and incorporate chessboard images within the scenes to facilitate camera calibration and pose estimation.
The dynamics in our dataset are notably more challenging than those in the DyNeRF dataset, featuring significant non-rigid motions of human bodies and transient effects such as slide effects on the screens in the backgrounds. To comprehensively assess the scalability of our method, we provide 5 scenes with a duration exceeding 200 seconds (6000 frames), which is over 10 times longer than the previous DyNeRF dataset.
This dataset was meticulously collected to evaluate the scalability of our method and to support further research. We have carefully reviewed the videos to ensure they do not contain any controversial or offensive content before using them for research purposes. While the faces in the videos may contain identifiable information, this information will be used solely for research and will not be disclosed to any third party. All videos were collected with the explicit consent of all participants. The view in the upper center is designated for validation.
@@ 3.2 Implementation details
Our implementation is based on HexPlane [7], Instant-NGP [31], and NGP-pl [25]. All the metrics are measured on a single RTX 3090 GPU if not otherwise specified. Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets. Following most of the state-of-the-art methods [28, 27, 17, 45, 43] we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset. We use $1280\times 720$ resolution for our dataset to fit into memory. We use the importance sampling scheme in [28] at a batch size of 1024 rays. The parameters are optimized by Adam [23], with $\epsilon=10^{-15},\beta=(0.9,0.96)$. The learning rate is set as 0.001 and scheduled by cosine function following [25]. We set the size of the hash table in the first branch as $2^{19}$ and the size of the hash table in the subsequent branches as $2^{14}$, and we set the structural parameters $(L,F)$ in the spatial hash encoders as $(12,2)$ if not otherwise specified. Following MERF [38] for static scenes, we use a smaller hash table for the hybrid representation in Eq. 6 than pure voxel hash grid [31] representation in Table 3 to ensure the compactness and scalability of our method. Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the CD-MERF are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively. The number of the parameters of the MERF representation is thus reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids. Since the hash tables take up the majority of the parameters, the model size is reduced from $21M$ to $7.5M$ as in Table 3.
At the base branch, we initialize the models and train for 18k iterations, and for each of the subsequent branches we train for 3k iterations using the previous branch for initialization. We use LeakyReLU activation in the MLP regressors. For long video sequences, we assign 30 chunks as an episiode. The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes. Strategies for updating the occupancy grids are identical with [31]. To save storage, we use only one shared occupancy grid for all the branches. We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision.
@@ 3.3 Results and comparisons
Table 1: Comparisons between our CD-NGP and SOTA methods on the DyNeRF dataset. Memory denotes the maximum host memory required for training. - denotes the metrics unreported in the original publications. Best, Second, and Third metrics are in highlight. The full model of [27] results in much worse metrics than those reported in their paper, so we only report the metrics of their pilot model.
@@
Table 1 shows the quantitative comparisons between CD-NGP and other state-of-the-art methods on DyNeRF dataset. We use the officially released metrics. The methods in [28, 34] are trained on 8 V100 GPUs or equally 4 A100 GPUs. The methods in [28, 34, 43] have no official implementations available. We are unable to obtain the missing quality metrics in [17] because of their huge memory occupation in the default resolution. HexPlane [7] uses resolution $(1024,768)$ in their implementation, so the memory occupation is much smaller than other offline methods. We use the official implementation of StreamRF [27] and INV [48] and report the average metrics on the 5 synchronized scenes of the dataset. StreamRF [27] and INV [48] are trained per-frame. NeRFplayer [43] is trained offline but streamable upon inference. 4DGS [55] caches the frames onto disk for training to alleviate memory occupation with a trade-off on training speed. The reconstruction quality is measured by PSNR, DSSIM, and LPIPS [57] (by AlexNet [24]) following common practices [28, 7, 45]. To prove the viability of our parameter-isolation-based method, we apply the previous state-of-the-art method HexPlane [7] to continual learning setting with parameter-regularization (HexPlane-PR) and generative replay (HexPlane-GR) as additional baselines.
@@
Table 2 and Fig. 5 shows the quantitative and qualitative comparisons on the meetroom dataset. Fig. 6 is the PSNR curve of the compared methods on the trimming scene. As shown in Table 2, HexPlane-GR and HexPlane-PR fail and produce unsatisfactory results on the meetroom dataset, we thus exclude them from Fig. 6. Our method outperforms the implicit MLP-based online method [48] in terms of quality, speed, and model size significantly. Moreover, our method consumes only a fraction of storage than the explicit voxel-based online baseline [27] while maintaining comparable reconstruction quality as compared in Fig. 5. When compared with offline baselines, our method produces comparable quality to MixVoxels [45] and outperforms HexPlane [7] in terms of quality and speed, and consumes much less memory as in Fig. 5. We use the released configuration of StreamRF [27] for the meetroom dataset. The full model of [27] results in much worse metrics than those reported in their paper, so we only report the metrics of their pilot model.
@@
CD-NGP (Ours) $\checkmark$ 26.01 0.039 0.1906 77min 113MB
3.3.3 Spatial representations.
[…]
CD-NGP 21M 30.23 0.198 75min 113MB
3D+4D 23M 29.34 0.243 82min 122MB
CD-NGP-C 21M 30.52 0.174 80min 114MB
CD-Plane 14M 29.16 0.215 77min 83MB
CD-MERF 7.5M 29.93 0.201 109min 57MB
CD-MERF-C 7.5M 30.22 0.174 115min 57MB
Our proposed continual learning framework is compatible with multiple spatial backbones. We provide quantitative results on hybrid representation (denoted as CD-MERF) and plane representation like [17, 21] (denoted as CD-Plane) on the DyNeRF dataset in Table 3. Since the element-wise product for plane representation in [17] produces unsatisfactory results due to the hash collision, we remove the $P_{m}(x,y,z)$ in Eq. 6 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation. We use $2^{P-2}$-length hash tables for the Plane representation in Table 3 to balance the quality and the model size. The structural configurations $(L,F)$ of the plane representation in Table 3 are set as $(6,4)$ for better results. The 3D+4D representation similar to [34] is yielded by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq. 5. By comparing our CD-NGP with the naive 3D+4D representation, we show the advantage of our asymmetric encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs. 3 and 7: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
Table 3 also shows metrics of different feature fusion methods. We replace the element-wise sum in Eqs. 7 and 6 with vector concatenation and denote them as CD-NGP-C and CD-MERF-C in Table 3. The hybrid representations (CD-MERF and CD-MERF-C) achieve comparable model sizes ($\sim$60MB) to methods for static scenarios [31, 21] without significant quality loss, reducing the average bandwidth cost to $0.2$MB/frame. Benefiting from larger feature dimensions, the vector concatenation leads to slightly better quality (especially when measured in LPIPS) by more training time, and the reasonable trade-off also shows the robustness of our method. Compared with the hash table designs in Table 7, the CD-NGP-C representation in Table 3 almost performs as well as the bulky $(18,18)$ configuration in Table 7 and reduces the model size by $80\%$. We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Table 4.
Model Flame Salmon Cook Spinach
[…]
CD-NGP 26.37 0.04586 0.2886 30.76 0.02440 0.1793
CD-NGP-C 26.65 0.04243 0.2680 31.03 0.02288 0.1593
CD-Plane 25.48 0.06058 0.3173 29.41 0.03487 0.2010
CD-MERF 26.15 0.04952 0.2921 30.35 0.02931 0.1858
CD-MERF-C 26.56 0.04435 0.2546 30.69 0.02609 0.1583
Model Flame Steak Sear Steak
[…]
CD-NGP 31.27 0.02132 0.1736 31.71 0.01917 0.1701
CD-NGP-C 31.64 0.01823 0.1403 32.02 0.01698 0.1424
CD-Plane 30.26 0.02949 0.1804 30.56 0.02723 0.1811
CD-MERF 31.01 0.02508 0.1719 31.44 0.02181 0.1698
CD-MERF-C 31.13 0.02424 0.1506 31.66 0.01989 0.1413
Model Cut Roasted Beef Average
[…]
CD-NGP 31.03 0.02366 0.1801 30.228 0.026882 0.19834
CD-NGP-C 31.26 0.02205 0.1578 30.520 0.024514 0.17356
CD-Plane 30.11 0.03302 0.1951 29.164 0.037038 0.21498
CD-MERF 30.71 0.02765 0.1856 29.932 0.030674 0.20104
CD-MERF-C 31.04 0.02450 0.1632 30.216 0.027814 0.17360
The detailed parameter numbers and bandwidth consumption are listed in Table 5. 3D Enc. and 2D Enc. denote the spatial encodings along voxels and planes respectively. The tuples $(M_{1},M_{2})$ in the cells denote the parameter numbers of the base encoder and each of the auxiliary encoders. Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth consumption $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$. Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
CD-NGP $(9.9M,0.37M)$ $0.042M$ $0.39M$ $21M$ 0.16 0.38
CD-NGP-C $(9.9M,0.37M)$ $0.046M$ $0.40M$ $21M$ 0.16 0.38
CD-Plane - $(1.9M,0.082M)$ $0.048M$ $0.28M$ $14M$ 0.11 0.28
CD-MERF $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.042M$ $0.16M$ $7.5M$ 0.061 0.19
CD-MERF-C $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.052M$ $0.17M$ $7.5M$ 0.065 0.19
@@
We also provide quantitative comparisons on our proposed long multi-view video dataset in Table 6 and Fig. 7 to further show the superior scalability of our method. We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset. We set the hash table size as $2^{12}$ in the auxiliary branches and use $T_{chunk}=5$ in our model. As shown in Table 6, our method reduces the memory and storage consumption by at least 4×4$\times$4 × with comparable quality to HexPlane [7]. Compared with the results in Table 1, the performance gap in storage consumption is even larger in long video sequences. This is because the whole model shares one base branch and the following auxiliary branches occupy the majority of the parameters. We thus set the hash table size in the auxiliary branches to $2^{12}LF$ and leverage the re-use of the features from the base branch to reduce the model size to $\mathcal{O}(N_{chunk}\cdot 2^{12}LF)$. .
The PSNR curves in Fig. 7 further demonstrate the stability of our method. The compared online methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than that in the end. The problem is properly eliminated by our method through parameter isolation, and our method continuously produce high-quality results for the whole sequence. We also provide comparisons between the rendered results in Fig. 8. The regularization method displayed in Fig. 8(a) reconstructs the TV screen in the background erroneously. The replay method displayed in Fig. 8(b) fails to recover the details. Ours in Fig. 8(f) produces comparable quality to the offline baseline in Fig. 8(c).
To further demonstrate the scalability of our method, we conduct experiments on the first 2400 frames of the seminar scene in our long video dataset, show the PSNR curves in Fig. 10 and compare the model sizes in Fig. 11. For simplicity of the evaluation, we set the chunk size as 10 for the CD-NGP and CD-MERF method. Our method produces high-quality results, and consumes only a fraction of storage space compared with the other online baselines.
Figure 7: PSNR curves of the first 1200 frames of seminar scene in our long video dataset. The solid lines and dotted lines denote offline and online methods respectively.